# Composed Multi-modal Retrieval: A Survey of Approaches and Applications

### v1

@@ 3.2.1 Design of Visual-Linguistic Combiner
Based on the key highlights of different combiner methods, ranging from simple to complex design structures, we classify them into the following 6 categories: (1) Global-level combination based on coarse-grained modality features; (2) Local-level combination focusing on fine-grained modality features; (3) Hierarchical combination that incorporates multiple feature granularities; (4) Combination that models structured relationship information between modality samples; (5) Combination that leverages external knowledge for enhancement; (6) Combination that considers data polysemy by adopting ambiguity probability modeling. Below, we introduce the details of each category.
Global-level Multimodal Combiner: Various methods have been proposed to effectively compose the global-level visual reference and textual modification into new semantic representations that resemble the visual target. The popular approach is based on gating and residual connection mechanisms. Vo et al. [18] introduce the Text Image Residual Gating (TIRG) method, which employs a gating mechanism and residual connection to modify image features based on text input while retaining the original image features when text modifications are insignificant. The gating mechanism computes a weighted sum of the gate and residual features, where the gate feature preserves the query image, and the residual feature introduces text-specific modifications. Dodds et al. [17] introduce a Residual Attention Fusion (RAF) mechanism for text-guided fashion image retrieval. RAF calculates attention weights from text features, applies them element-wise to image features, and uses residual connections to maintain alignment in the embedding space. This approach achieves global modifications while preserving the primary structure of the image features. Baldrati et al. [19] design a Combiner network that maps image and text features into a shared embedding space using a pre-trained CLIP model. The Combiner processes concatenated features through linear layers, ReLU activations, and Dropout to generate a combined feature. In the subsequent works [20, 21], they fine-tune CLIP’s text encoder for task-specific adaptation and propose a Combiner that outputs a normalized combination of text and image features, including convex and learned mixtures. Lin et al. [11] also fine-tune CLIP’s encoders and employ a fusion module with weighted summation, concatenation, and bilinear pooling to capture complex modification intents.
Besides, Lee et al. [22] present the Content-Style Modulation (CoSMo) framework, which decomposes textual features into content and style components. Content modulates spatial activation patterns of image features, while style adjusts their distribution via learnable affine transformations. An adaptive fusion module dynamically balances these components based on the text, optimizing multimodal feature integration. Anwaar et al. [23] design a method based on complex space mapping and symmetry constraints. Specifically, they map image and text features into a complex space. Text features define rotational transformations from source to target image features, with a rotational symmetry constraint, ensuring that multiplying the target image feature with the complex conjugate of the query text feature yields a representation similar to the query image feature. Deep metric learning is used to optimize the composition and mapping of these complex vectors. Xu et al. [24] introduce a composition-and-decomposition framework. The composition network couples reference image and text features into a joint representation, while the decomposition network decouples the target image into visual and textual subspaces. An attention mechanism selects key components, and techniques like bilinear pooling enhance feature extraction, forming a closed-loop system for mutual optimization.
Local-level Multimodal Combiner: Current local feature-based CMR methods can be primarily categorized into three approaches: attention-based feature interaction methods [25, 26, 27] that achieve fine-grained correspondence through multi-layer or bidirectional attention mechanisms; semantic enhancement and suppression methods [28, 29, 30] that focus on selective modulation of specific semantics via dynamic weights or kernels; and concept alignment and distribution learning methods [31, 32] that emphasize visual-semantic alignment at the conceptual level and optimize matching boundaries. While these approaches differ in their technical strategies, they all aim to improve the utilization of local features for more precise retrieval performance.
Attention-based feature interaction methods focus on establishing fine-grained correspondences between image regions and text words. For instance, Hosseinzadeh and Wang [25] employ a multi-layer bidirectional attention mechanism to enhance the understanding of specific image region modifications in relation to text instructions. Each layer’s two sub-modules compute linguistically attended visual features and visually attended linguistic features, creating a detailed interaction between the visual and textual features and gradually generating richer joint representations. Similarly, FashionViL by Han et al. [26] introduces a dynamic, learnable attention module to build region-word correspondences, further refining feature fusion through bilinear pooling, which captures second-order statistics to better integrate complex visual and textual features. Zhang et al. [27] also propose to achieves deep integration of local modal features through Multi-modal Complementary Fusion (MCF) and Cross-modal Guided Pooling (CGP) modules. The MCF module uses a double-interaction mechanism to compute attention weights and interaction maps between the reference image and text, projecting text features into the visual space and fusing them with image features via a modified gating operation to enrich the image representation with key textual information. The CGP module explores both intra-modality and inter-modality relationships, calculating cross-modal relevance between local image features and text word vectors, as well as intra-modal interactions among channels of the reference image. It dynamically assigns weights to local features, generating a holistic representation that integrates cross-modal information. This design emphasizes leveraging local modal features for finer-grained interactions and more accurate retrieval outcomes.
Semantic enhancement and suppression methods, on the other hand, center around selectively enhancing or suppressing specific semantic features of the image based on textual information. Pang et al.’s Weak-text Semantic Augment (WSA) module [28] enhances short or incomplete text queries by incorporating relevant image information. WSA expands the text feature dimensions and uses a multi-layer perceptron to map image features into the text’s latent space. An attention mechanism then integrates these image features and maps them into the semantic space to enhance the text’s semantic expression. This allows WSA to flexibly select high-level image features pertinent to the text, improving the understanding of user query intent, especially for brief or incomplete text descriptions, thus enhancing retrieval accuracy and robustness. Zhang et al.’s approach [29], involving multi-semantic dynamic suppression and text-semantic complementary selection, divides the retrieval task into two sub-processes: suppressing irrelevant image details and enhancing relevant ones based on the text, ensuring that the final image representation aligns closely with the textual description. Similarly, the SAC framework [30] proposed by Jandial et al. uses a two-stage approach, first identifying salient regions in the image relevant to the text and then adjusting these regions while preserving other areas of the image, employing a combination of LSTM and batch normalization for feature fusion. These methods are united by their focus on the selective enhancement or suppression of semantic features, using dynamic weights or kernels to modulate feature importance while maintaining the integrity of unmodified areas.
Finally, concept alignment and distribution learning methods emphasize aligning visual and textual concepts at the semantic level. Zhao et al. [31] achieve concept-level visual-semantic alignment. By concatenating feature tokens of the reference image and target image, and then applying a Transformer to perform context modeling on the concatenated features, the model subsequently employs an attention-based multiple-instance learning framework to compute alignment scores between visual concept features and semantic concept embeddings in the text modifiers. This alignment is then used to guide the fusion of the reference image and text features, improving retrieval accuracy. Similarly, Wang et al. [32] explore the positive and negative correspondences of local features, introducing a threshold optimization mechanism to dynamically distinguish between matched and unmatched samples. By optimizing this threshold, the module strengthens the model’s ability to distinguish between local relevances and irrelevances, enabling it to better handle highly similar samples and avoid mismatches.
Hierarchical Multimodal Combiner: The design of hierarchical multi-modal combiner is driven by the need to capture multi-granular visual and textual relationships, enabling a more comprehensive fusion of information across different levels of abstraction. Existing methods can be categorized into two primary types based on how hierarchical features are generated: encoder multi-layer strategies [33, 34, 35, 36, 37, 38, 39, 40] and global-local feature strategies [41, 42, 43, 44, 29].
The first type approaches utilize features extracted from different layers of encoders, such as CNNs or Transformers, allowing for a more detailed representation of image-text interactions. Chen et al. [33] propose a method where visual features from multiple layers of a CNN are concatenated with textual features, forming a visiolinguistic representation. This representation is then processed through two attention mechanisms: a self-attention stream that uncovers region-to-region dependencies and a joint-attention stream that modulates the preservation of visual content based on the textual input. The use of features from different CNN layers enables the model to preserve both high-level semantics and low-level visual details, enhancing the overall multigranular fusion of information. Dodds et al. [34] propose a multimodal retrieval model that extracts image features from different ResNet50 layers, treating them as tokens representing various parts of the image. These tokens, along with tokens generated from textual features, are concatenated into a unified sequence. Through self-attention, the model enables interaction between image and text tokens, facilitating cross-modal fusion. Using features from different ResNet layers, the model captures both global and local visual characteristics, making it effective for retrieving target concepts from modified query images.
Similarly, Jandial et al. [35] present a framework that extracts multi-scale visual features from a CNN, forming a visual feature pyramid. Each level of this pyramid is aligned with textual embeddings generated by a GRU, and these multimodal embeddings are further processed through a Semantic Feature Transformation module. This module uses attention mechanisms to align the visual and textual information at each level, followed by a Hierarchical Feature Aggregation module that aggregates these multi-level representations. This hierarchical processing ensures a comprehensive visiolinguistic composition, allowing the model to address the complexities of text-conditioned image retrieval. Yang et al. [36] introduce a model that integrates modification and auxiliary relation modeling, using a multi-level aggregation method to fuse visual and textual information. Through the Hierarchical Aggregation Transformer, the model captures both low-level texture details and high-level semantic information, ensuring a richer representation for multi-modal composition. This design enhances the model’s flexibility in adapting to different types of query texts and improves the retrieval process by ensuring a more accurate alignment between reference images and target images.
These methods leverage hierarchical feature extraction to further localize the modified specific image regions. Zhang et al. [37] focus on image localization for composed image retrieval, employing a multi-level approach that leverages feature maps from shallow, middle, and deep layers of the CNN. The localization process is applied across these layers, using up-sampling and down-sampling operations to align the localization distribution, resulting in an improved fusion of multi-modal features. This approach enhances the accuracy of localizing relevant areas of the image in response to textual input, thereby supporting precise multi-modal feature integration. A similar multi-level design is evident in Xu et al.’s approach [38], where hierarchical feature maps from a Vision Transformer are used to capture visual details at various levels. Then, the designed global-local alignment bridges the gap between composed queries and target images by focusing on both global embedding spaces and local region details through spatial attention mechanisms. This design improves the model’s ability to refine specific areas of the reference image according to the modification text, enhancing retrieval accuracy. Li et al. [39] extend this idea with a method that extracts multi-grained visual and textual features from different layers of a pre-trained ResNet50. By incorporating a Multi-Grained Semantic Construction module, they capture fine, mid, and coarse-grained information from both modalities. Their attention mechanism, supervised by a triplet ranking loss, ensures that textual features focus on the modified regions while excluding the preserved ones, optimizing the multi-modal fusion process. Besides, Huang et al. [40] introduce the language-guided local infiltration system, which focuses on hierarchical feature extraction and fusion to modify specific image regions while preserving the overall structure. Specifically, The language prompt visual localization module generates a localization mask to identify the target region for modification based on the text, enabling precise control over which part of the image is altered. The text infiltration with local-awareness module then applies the mask to adjust the image, ensuring that local changes align with the text while maintaining the integrity of other regions.
The second type approaches emphasize a hierarchical approach to feature extraction, where local, regional, and global information is progressively integrated, ensuring a rich representation for multimodal retrieval tasks. Wen et al. [41] propose a method that incorporates both local and global composition modules to handle multimodal queries. The local composition module focuses on fine-grained visual descriptors, while the global composition module integrates global visual features with textual modifications. These modules use attention mechanisms to align the image and text, allowing the model to capture detailed local interactions as well as broader semantic shifts. A mutual enhancement module is introduced to facilitate knowledge sharing between the two composition modules, reinforcing the performance of both. This hierarchical design enables the model to handle a wide range of modification types, from specific local changes to more abstract visual adjustments, by combining these different levels of information. Liu et al. [42] introduce a model that leverages a pre-trained vision-and-language (V&L) model, OSCAR, to map composed queries consisting of a reference image and textual description into an embedding space. While OSCAR processes local text sequences, the model adapts to images by using a pre-trained ResNet to extract global image features. This design allows for the combination of global image features with modification text, processed through a multi-layer transformer structure with self-attention mechanisms. By using global image features instead of region-specific representations, the model captures a broader understanding of the image and text relationship, while still leveraging local text sequence processing.
Goenka et al. [43] introduce a fashion image retrieval model that extracts features at multiple levels: the global image, cropped clothing regions, regions around fashion landmarks, and areas identified by a pre-trained object detector (RoIs). These features are then integrated with textual feedback through a Vision-Language Transformer, while the target image processing focuses solely on visual features to reduce inference costs. This method enhances the fusion of visual cues and textual information, facilitating efficient retrieval. Wen et al. [44] design a multi-granularity feature processing mechanism using fine-grained local and global composition modules. Their model assigns specific matching tokens for attributes like color, belt, and collar, which adaptively aggregate features during training. This approach enables the effective use of multimodal data, from fine-grained details to coarser, more abstract information. Zhang et al. [29] integrate CNNs with Transformers for image feature extraction, where the CNN captures local features while the Transformer models global edge information. This hybrid approach improves the representation of reference images by combining local and global features, while LSTMs handle textual feature extraction. The final fusion of these features through residual and gating mechanisms enhances the model’s ability to match query-image and target-image pairs.
In summary, these methods emphasize hierarchical feature extraction from various encoder layers to effectively capture both low-level and high-level visual and textual information. By incorporating attention mechanisms and leveraging different feature granularities, these methods improve the flexibility and precision of multi-modal composition, leading to more accurate retrieval and understanding of composed queries.
(4) Structured relational multimodal combiner: Some studies collectively advance the field of multi-modal composed retrieval, from foundational hierarchical composition learning to innovative applications of residual learning, comprehensive relationship reasoning, complementary association learning, and style-guided patch interaction. While these research methods focus on different technical strategies, they all aim to enhance the compositional modeling of multi-modal information from various perspectives, improving the retrieval performance of complex composed queries.
Xu et al. [38] introduced a Hierarchical Composition Learning (HCL) framework to tackle hierarchical composition problems. This method progressively integrates complex relationships between text and images through a multi-level compositional structure. Similarly, Li et al. proposed the Residual Learning for Text and Image Composition using Graph Convolutional Network (RTIC) method [45], introducing residual connections to enhance information flow between text and images and improve retrieval accuracy and robustness. RTIC incorporates a residual learning module based on GCNs, which accumulates and propagates differences between modalities through multiple residual connections, and an adaptive weighting mechanism to dynamically adjust the importance of different layers.
Zhang et al. [46] developed comprehensive relationship reasoning, which utilizes explicit graph structures to capture intricate relations among elements in text and images. The approach dynamically constructs and updates graph structures to accommodate various querying needs. Moreover, Jiang et al. [47] presented Complementary Association Learning (CaLa), emphasizing the importance of mining complementary information from texts and images. CaLa employs a dual-stream encoder with cross-attention mechanisms to enhance retrieval performance. Most recently, Chen et al. introduced the Style-guided Patch Interaction (SPIRIT) method [48], addressing text-feedback-based image retrieval in the fashion domain by enhancing interactions between image patches through the incorporation of style information. SPIRIT utilizes a style encoder to extract style features from input images, employs a style-guided attention mechanism to adjust the interaction strength between image patches dynamically, and incorporates a text feedback module to further optimize retrieval results based on user-provided text input.
(5) Knowledge-enhanced multimodal combiner: Several approaches have been proposed to enhance the performance of multi-modal compositional retrieval by leveraging external knowledge or specific mechanisms. Among these, Zhang et al. [49] emphasizes the modeling of geometric structures and cross-modal relationships. Moreover, Zhang et al. [3] and Liu et al. [50] further improve upon this by introducing external knowledge and sophisticated negative sample generation strategies to optimize multi-modal fusion. In contrast, Bai et al. [51] focuses on utilizing sentence-level prompts to improve retrieval accuracy and generalization capability. Specifically, Zhang et al. [49] proposed a Geometry-Sensitive Cross-Modal Reasoning network (GSCR), which achieves joint modeling of the visual-semantic relationships and geometric structures of images with modified texts through a Geometry-Sensitive Inter-Modal Attention module (GS-IMA) and a Text-Guided Visual Reasoning module (TG-VR). While GS-IMA explicitly and implicitly incorporates spatial structure into cross-modal attention, TG-VR is utilized to model semantic information not present in the reference image, guiding further visual reasoning. This design emphasizes the importance of leveraging the spatial structures of images and enhancing cross-modal interactions when dealing with complex queries. Similarly, Zhang et al. [3] introduced Multi-modal Compositional Example Mining (MCEM), an approach that optimizes fusion across different modalities by generating component-level negative samples. By pairing logically mismatched sentences with reference images and introducing a novel sentence augmentation method to generate more challenging multi-modal negative samples, MCEM assists the model in learning a better metric space. This method highlights the importance of creating sophisticated negative samples to improve the understanding of multi-modal data by the model, showcasing innovation, particularly in utilizing external knowledge. Building upon this foundation, Liu et al. [50] presented a Dual Multi-modal Encoder-based Re-ranking method (DMER), which uses two independent multi-modal encoders to encode images and text separately, then re-ranks candidate images based on their similarity scores. By integrating an external knowledge graph, DMER enables the model to leverage additional semantic information for reasoning, thereby enhancing its expressiveness. This approach underscores the value of utilizing an external knowledge graph to supplement traditional single-encoder methods, demonstrating significant effectiveness especially on large-scale datasets. Bai et al. [51] introduced Sentence-Level Prompts Benefiting Composed Image Retrieval (SLPCIR), aiming to enhance CIR performance through sentence-level prompts. By parsing textual descriptions to identify key sentence constituents and generating corresponding prompts for each constituent, SLPCIR guides the model in more accurately locating and retrieving target images. This method illustrates how sentence-level prompts can be used to improve retrieval accuracy, showing strong generalization capabilities particularly in zero-shot and few-shot scenarios, thus enhancing the effectiveness of multi-modal compositional retrieval.
(6) Ambiguous uncertainty/probabilistic multimodal combiner: Neculai et al. [52] addressed the challenge of combining multiple query inputs in multimodal image retrieval by proposing a Multimodal Probabilistic Combiner (MPC). Existing approaches are typically limited to handling single or dual query inputs, which constrains their capability to cope with complex scenarios. The MPC is designed to retrieve target images based on an arbitrary number and types of queries. To this end, the authors established a new benchmark using the MS-COCO dataset to evaluate the performance of multimodal image retrieval under varying modalities and query counts. The MPC learns probabilistic embeddings and employs parameterized probabilistic combination rules to synthesize these embeddings, resulting in composite multivariate Gaussian distributions that capture both the semantics and uncertainty of the input data. Xu et al. [53] focused on addressing the issues of semantic polysemy and data uncertainty. Traditional one-to-one mapping methods struggle to effectively handle these challenges due to their inability to adequately capture the diversity of multimodal data. To tackle this issue, the authors introduced the Set of Diverse Queries with Uncertainty Regularization (SDQUR). SDQUR leverages the Q-Former from BLIP2 to extract multiple deterministic embeddings from unimodal or multimodal inputs, where each embedding represents a potential matching factor. To address data uncertainty caused by label noise, SDQUR incorporates an uncertainty regularization module that encodes the input as a distribution, allowing for the sampling of multiple instances to establish probabilistic many-to-many correspondences.

### v2

@@ 3.2.1 Design of Visual-Linguistic Combiner
As illustrated in Fig. 3, we classify existing combiner methods into the following 6 categories, ranging from simple to complex design structures: (1) Global-level combination based on coarse-grained modality features; (2) Local-level combination focusing on fine-grained modality features; (3) Hierarchical combination that incorporates multiple feature granularities; (4) Combination that models structured relationship information between modality samples; (5) Combination that leverages external knowledge for enhancement; (6) Combination that considers data polysemy by adopting ambiguity probability modeling. Each combiner requires to solve two issues: the first is ”where to see”, i.e., to align the semantics between the modified text and the reference vision to find the content that needs to be modified, and the second is ”how to change”, i.e., to fuse multimodal information to modify the visual and language semantics, conveying more flexible retrieval intentions. Therefore, in the following, we will introduce the details of cross-modal alignment and fusion for each type of combiner.
(1) Global-level Multimodal Combiner: Global-level alignment occurs through interactions between image and text modalities exclusively at their global feature representations within a shared semantic space. Recent dual-stream models, including ViLBERT [159], UniVL[160], ERNIE-ViL [161], have received more attention. They utilize separate visual and textual encoders to project images and texts into a shared space for semantic similarity measuring. The most representative work is CLIP [162], which has demonstrated remarkable capabilities. CLIP employs dedicated visual and textual encoders, learning semantic associations through contrastive optimization on 400 million image-text pairs.
In the CMR field, various methods have been proposed to effectively fuse global-level visual references and textual modifications into unified semantic representations, as illustrated in Fig. 3(a). Early approaches focused on adaptive feature integration mechanisms. Vo et al. [20] integrate gating mechanisms with residual connections, computing weighted combinations of original image representations and text-induced modifications to preserve visual content while enabling semantic adjustments. Dodds et al. [32] employ residual attention fusion, deriving attention weights from textual features and applying them element-wise to image features for global semantic adaptation. Recent CLIP-based methods have gained prominence for their effectiveness. Baldrati et al. [33, 34, 35] design Combiner networks that process concatenated CLIP features through linear layers and activations, later extending to normalized combinations including convex and learned mixtures. Lin et al. [36] fine-tune CLIP encoders with fusion modules employing weighted summation, concatenation, and bilinear pooling to capture complex modification intents. In addition, advanced composition strategies explore specialized feature modeling approaches. Lee et al. [37] propose CoSMo, disentangling textual features into content and style components that modulate spatial activation and distribution of image features through adaptive fusion. Anwaar et al. [38] develop complex-space transformations where text features apply rotational operations to image features under symmetry constraints. Xu et al. [39] introduce composition-decomposition frameworks with joint encoding and target image decoupling, employing attention mechanisms and bilinear pooling in closed-loop structures for enhanced feature interaction.
(2) Local-level Multimodal Combiner: Local-level feature alignment necessitates the capture of fine-grained visual-textual correspondences to enable precise cross-modal semantic measurement at the fragment level. Pioneering work, by Karpathy and Fei-Fei [163], first attempts to optimize the most similar region-word pairs for selecting matched semantics. The Stacked Cross Attention Network (SCAN) [9], as a representative work, employs any fragment from one modality as a query to interact with all fragments from the complementary modality. SCAN inspires many variants, such as focal attention [164, li2024improving], iterative attention [165], and negative-aware attention [166]. In recent years, pre-trained models based on the Transformer attention mechanism for fine-grained semantic alignment have also received great attention. This type of model integrates different modal information at an early stage, and because all modal information is interactively processed through only one branch, this type of model is also called a single-stream model. Representative works include Unicoder-VL [167], ImageBERT [168], UNITER [169], BLIP2 [170], etc.
Current local feature-based CMR methods can be primarily categorized into two approaches: attention-based feature interaction methods [40, 41, 42] that achieve fine-grained correspondence through cross-modal attention mechanisms; semantic enhancement and suppression methods [43, 44, 45, 46, 47] that focus on selective modulation of specific semantics via dynamic weights or kernels. While these approaches differ in their technical strategies, they all aim to improve the utilization of local features for more precise retrieval performance. Attention-based feature interaction methods aim to construct fine-grained correspondences between visual and textual modalities by employing various forms of attention mechanisms. Hosseinzadeh and Wang [40] utilize a multi-layer bidirectional attention framework in which each layer computes both linguistically attended visual features and visually attended linguistic features, progressively enriching the joint representation. Han et al. [41] develop a dynamic, learnable attention module to establish region-word alignments and apply bilinear pooling to capture second-order interactions, thereby enhancing feature integration. Zhang et al. [42] introduce Multi-modal Complementary Fusion and Cross-modal Guided Pooling, combining dual-attention-based projection of text into visual space, gated fusion, and modality-aware pooling to dynamically weight local features. These methods share the objective of improving cross-modal feature interaction through layered or dual-attention structures and enhanced fusion mechanisms.
Semantic enhancement and suppression methods, on the other hand, center around selectively enhancing or suppressing specific semantic features of the image based on textual information. Pang et al. [43] integrate relevant image features into an expanded text representation via a multi-layer perceptron and attention, enhancing understanding of short or incomplete queries. Zhang et al. [44] suppress irrelevant image details while enhancing relevant ones through dynamic semantic weighting to align images with text better. Jandial et al. [45] first detect salient image regions linked to text, then adjust these while preserving other areas, using LSTM and normalization for feature fusion. These methods are united by their focus on the selective enhancement or suppression of semantic features, using dynamic weights or kernels to modulate feature importance while maintaining the integrity of unmodified areas. Extending this line of research, Zhao et al. [46] propose a concept-level visual-semantic alignment framework by concatenating feature tokens from both reference and target images and employing a Transformer for joint context modeling. An attention-based multiple-instance learning module then calculates alignment scores between visual concept representations and textual semantic embeddings, guiding the fusion process to ensure concept-level consistency in retrieval tasks. Wang et al. [47] introduce a threshold optimization strategy to refine the identification of positive and negative local feature correspondences. By dynamically adjusting decision thresholds, the model improves its capacity to discriminate between visually similar but semantically distinct instances, thereby reducing retrieval errors.
(3) Hierarchical Multimodal Combiner: This design is driven by the need to capture multi-granular visual and textual relationships, enabling a more comprehensive fusion of information across different levels of features. The alignment of semantic features of different granularities in visual language has been studied for a long time [171, 172, 173]. Early work proposed to model the semantic alignment of global and local features of vision and language [174, 175]. On this basis, to progressively mine the semantic alignment relationship between modalities, Chen et al. [165] proposed a serialized multi-level alignment model with iterative operation, Hu et al. [176] employ a multi-layer CNN to capture the local correlations and long-term dependencies between vision and language, and Ji et al. [177] proposed a local-global cross-interaction alignment model.
Existing hierarchical multimodal combiners can be categorized into two primary types based on how hierarchical features are generated: encoder multi-layer strategies [48, 49, 50, 51, 124, 52, 53, 54] and global-local feature strategies [55, 21, 56, 57, 44]. The first type of approaches utilizes features extracted from different layers of encoders, such as CNNs or Transformers, allowing for a more detailed representation of image-text interactions. Several works [48, 49, 50, 51] utilize features from different CNN or Transformer layers, which are fused with textual features using attention mechanisms. For example, Chen et al. [48] apply both self-attention and joint-attention to capture intra-modal and cross-modal relations. Similarly, Dodds et al. [49] treat intermediate CNN features as visual tokens and align them with text via transformer-based attention, facilitating multi-granular fusion. Others [50, 51] explicitly construct multi-level visual pyramids aligned with text embeddings, which are then aggregated through specialized modules such as semantic feature transformation or hierarchical transformers. These designs support alignment across visual scales and improve visiolinguistic composition. A subset of methods [124, 52, 53, 54] extend the above designs by incorporating spatial localization mechanisms. For instance, Zhang et al. [124] perform multi-layer localization alignment, while Xu et al. [52] combine hierarchical vision features with global-local spatial alignment. Li et al. [53] construct multi-grained representations for focused region modification, and Huang et al. [54] employ language-guided masks to direct local feature adjustment. These approaches aim to better ground textual modifications in corresponding image regions.
The second type emphasizes a hierarchical approach to feature extraction, where local, regional, and global information is progressively integrated, ensuring a rich representation for multimodal retrieval tasks. Several methods [55, 57] employ distinct local and global composition modules, where fine-grained attribute alignment is combined with holistic visual-textual fusion, often facilitated by mutual learning components. Liu et al. [21] utilize pre-trained vision-and-language models to encode local text and global image features, followed by transformer-based attention fusion, enabling abstract and broad semantic understanding. In more specialized domains, models such as [56] extract visual features at multiple spatial levels, including full images, object-centric crops, and region-of-interest segments, which are fused with textual feedback through vision-language transformers. Hybrid architectures [44] integrate convolutional and transformer networks to jointly capture local textures and global scene structure, followed by feature fusion via residual or gating mechanisms for robust cross-modal matching.
(4) Structured Relational Multimodal Combiner: Recent methods have significantly advanced composed multimodal retrieval by explicitly capturing and utilizing structural dependencies between visual and textual inputs. Leveraging structural dependencies can promote more accurate cross-modal semantic alignment. Early work introduced visual and linguistic scene graphs to enable context-aware alignment [178, 179, 180, 181, 182], learning modality-specific relations independently. A representative method, SGM by Wang et al. [178], designs visual and textual scene graph encoders that enrich node representations through neighborhood aggregation, thereby extracting cross-modal features at object- and relation-level for alignment interaction. Moving beyond reliance on predefined scene graph extractors, the adoption of graph convolutional networks (GCNs) [183, 184] for non-Euclidean spatial data has enabled more flexible graph-based alignment strategies [185, 186, 187, 188, 189, 190, 191, 192]. Liu et al. [185] propose to model the graph structures of images and texts respectively, and use GCNs to construct fine-grained phrase-level alignment. Inspired by this, Diao et al. and Zhang et al. enhanced graph structure alignment from the perspectives of filtering and confidence [188, 189], respectively, while works [193, 194] propose hierarchical graph structure interactions. In contrast to separate modeling modality graphs, multimodal graph structures [195, 196, 197] that jointly represent both visual and textual elements have also been proposed, offering a unified framework to strengthen cross-modal interaction and fusion.
The structured relational multimodal combiner collectively emphasizes the importance of modeling hierarchical, complementary, and context-dependent relationships. A line of work focuses on modeling visual and textual graphs [125]. Jiang et al. [60] introduce a dual-stream architecture with cross-attention to mine and align complementary information between modalities. Further emphasizing context-dependent relationships, Chen et al. [61] incorporate structured style modeling to guide intra-image patch interactions in a text-conditioned manner, refining patch-level relationships. Xu et al. [52] propose a hierarchical composition framework that progressively models visual graphs with entity, attribute, and relationship nodes, facilitating the gradual fusion of complex visual-textual associations. Li et al. [58] enhance the flow of cross-modal information through residual learning with GCNs, effectively preserving and propagating modality-specific differences through structured residual connections. Zhang et al. [59] employ dynamic graph construction to model detailed semantic relationships between multimodal elements, supporting more flexible and context-sensitive retrieval.
Another line of work proposes to model the unified multimodal graph. Zhang et al. [126] construct a unified relational graph, where nodes represent visual elements enriched with textual attributes, and edges capture relationships among these attributes. Nie et al. [127] propose a multimodal hierarchical graph neural network to effectively model conversational structures and multimodal contexts. They employ a graph attention network to dynamically aggregate information from different modalities for accurate user intent understanding. Zhang et al. [62] incorporate textual semantics into visual nodes, enabling each node to embed both visual features and corresponding textual information for enhanced multimodal representation.
(5) Knowledge-Enhanced Multimodal Combiner: By incorporating external knowledge, these approaches aim to enrich the semantic representation space, reinforce cross-modal alignment. Three types of knowledge are usually introduced: geometric information, dependency relations, and common sense knowledge. Regional geometric information [198], obtained by object detectors [199, 200], is typically used to enhance visual semantic features. Among them, Wang et al. [201] propose an attention network based on position features, Liu et al. and Zhang et al. respectively propose geometric graph connection methods based on the relative positions of image regions [185, 189]. Dependency relations are typically obtained by off-the-shelf toolkits such as Spice [202], Stanford CoreNLP [203], or visual scene graph generation tools such as MSDN [204] and NeuralMotifs [205]. Introducing structured dependency knowledge can significantly improve the alignment accuracy [206], as discussed in Sect. 3.2.1(4). Common-sense knowledge, including scene co-occurrence [207], corpus co-occurrence [208], and action information [209], is used to further enhance the accuracy of visual-language alignment.
Introducing external knowledge also improves the model’s capacity to interpret and reason over complex compositional queries. Zhang et al. [62] propose a geometry-sensitive cross-modal reasoning network that integrates visual-semantic and spatial structural information through an inter-modal attention mechanism and a visual reasoning module guided by text. This method allows for explicit modeling of spatial relationships while enabling semantic extrapolation beyond the visual content of reference images. Zhang et al.  [71] design a compositional example mining strategy that constructs challenging negative samples by pairing mismatched image-text pairs and augmenting them with compositional variants. This technique enhances the model’s ability to learn discriminative representations by refining the decision boundary in the joint embedding space. In a complementary direction, Bai et al. [63] propose a sentence-level prompting framework that decomposes textual descriptions into semantically meaningful constituents. These constituents are then used to generate structured prompts, which guide the model in locating corresponding visual content more precisely.
(6) Ambiguous Probabilistic Multimodal Combiner: As shown in Fig. 3(f), to address ambiguity and uncertainty problem, several approaches have been developed that incorporate probabilistic modeling. In vision-language alignment, a typical paradigm is to encode each input into a set of embeddings [210, 211, 212, 213, 214]. Song et al. [210] propose to regularize the learned embedding space by minimizing the discrepancy using the maximum mean discrepancy. Chun et al. [211] propose to use probabilistic embeddings to represent vision-language as probability distributions in a common embedding space. Kim et al. [212] propose a set prediction module and smooth-Chamfer similarity for set-based embeddings. Zhang et al. [214] propose a novel set-embeddings-based method, which improves accuracy and efficiency from the perspective of dissecting key semantic dimensions within each subspace.
In composed multimodal retrieval, Xu et al. [65] introduced the Set of Diverse Queries with Uncertainty Regularization framework to address challenges arising from semantic polysemy and noisy supervision. This method extracts multiple deterministic embeddings using a vision-language encoder and encodes them into a distributional form to model uncertainty. A regularization module is employed to enable sampling-based probabilistic matching, thus establishing robust many-to-many correspondences between multimodal inputs and retrieval targets. Neculai et al. [64] proposed the Multimodal Probabilistic Combiner, which extends the capacity of retrieval systems to handle an arbitrary number and types of query modalities.